OpenStudy (loser66):

Prove that \(2^{2^{n+1}} \equiv 1(mod~2^{2^n}} + 1\) Please, help

OpenStudy (loser66):

\(2^{2^{n+1}} \equiv 1 (mod~2^{2^n} +1)\)

jimthompson5910 (jim_thompson5910):

use the fact that a = b (mod m) ---> m*k = a-b to turn 2^(2^(n+1)) = 1 (mod 2^(2^n) + 1) into (2^(2^n) + 1)*k = 2^(2^(n+1)) - 1 if we can show k is an integer, then we've effectively proven the equation above is true (2^(2^n) + 1)*k = 2^(2^(n+1)) - 1 (2^(2^n) + 1)*k = 2^(2^n*2^1)) - 1 ... note1 (2^(2^n) + 1)*k = 2^(2^n*2) - 1 (2^(2^n) + 1)*k = [2^(2^n)]^2 - 1 ... note2 (2^(2^n) + 1)*k = [2^(2^n) - 1][2^(2^n) + 1] ... note3 k = 2^(2^n) - 1 ... note4 so this proves k is indeed an integer since 2^(2^n) is an integer note1: used the identity a^x*a^y = a^(x+y) note2: used the identity (a^x)^y = a^(x*y) note3: used the difference of squares factoring rule note4: divided both sides by 2^(2^n) + 1

jimthompson5910 (jim_thompson5910):

n is an integer 2^n is an integer (integer to any integer power = integer) 2^(2^n) is an integer 2^(2^n) - 1 is an integer (integer - integer = integer)

OpenStudy (loser66):

Got you. Thank you so much.

OpenStudy (loser66):

I need know the easiest way to solve \(3^{323} mod 17\). Please

jimthompson5910 (jim_thompson5910):

hints 323 = 17*19 https://en.wikipedia.org/wiki/Fermat%27s_little_theorem

jimthompson5910 (jim_thompson5910):

and 19 = 17+2 so 3^19 = 3^(17+2) = 3^17*3^2

OpenStudy (loser66):

got it, too. :)